OpenStudy (anonymous):

The altitude of an equilateral triangle is 10. What is the length of each side?

OpenStudy (bahrom7893):

a

OpenStudy (anonymous):

A. 20 sqrt 3 B. 10 sqrt 3 C. 20 sqrt 3/3 D. 10 sqrt 3/3

OpenStudy (bahrom7893):

using pythagorean thm: a^2=a^2/4+100

OpenStudy (bahrom7893):

a^2-a^2/4=100 3a^2/4=100 3a^2=400